Palomar Battery Company expects to operate at 75% of full capacity during April. The total manufacturing costs for April for the production of 60,000 batteries are budgeted as follows:

            Direct materials                        $ 75,000

            Direct labor                             960,000

Variable factory overhead       111,000

Fixed factory overhead           288,000

Total manufacturing costs      $1,434,000

The company has an opportunity to submit a bid for 17,500 batteries to be delivered by April 30 to a government agency. If the contract is obtained, it is anticipated that the additional activity will not interfere with normal production during April or increase the selling administrative experiences.

  1. What is the April budgeted cost per battery for the production of 60,000 batteries?’
  2. What is the unit cost below which Palomar Battery Company should not go in bidding on the government contract?
Answer #1

SOLUTION

(A) The April budgeted cost per battery for the production of 60,000 batteries = $23.9 per battery

Total Budgeted cost =Direct material + Direct labor+ Variable factory overhead+ Fixed factory overhead

= 75,000 + 960,000 + 111,000 + 288,000 = 1,434,000

Budgeted cost per battery of 60,000 batteries =Total budgeted cost / Total production

= $1,434,000 / 60,000 = $23.9

(B) The unit cost below which Palomar Battery Company should not go in bidding on the government contract is = $19.1 per unit

Total variable cost = Direct material + Direct labor+ Variable factory overhead

= 75,000 + 960,000 + 111,000 = 1,146,000

Variable cost per battery = 1,146,000 / 60,000 = $19.1